OpenStudy (anonymous):

Solve the equation : 21x2 -28x +10=0.

sam (.sam.):

21x^(2)-28x+10=0 x=(-b+- sqrt(b^(2)-4ac))/(2a) where ax^(2)+bx+c=0 x=(-(-28)\ sqrt((-28)^(2)-4(21)(10)))/(2(21)) x=(28+- sqrt((-28)^(2)-4(21)(10)))/(2(21)) x=(28+-2i sqrt(14))/2(21) x=(28+-2i sqrt(14))/(42) --------------------------------------- x=(14+i sqrt(14))/21 x=(14-i sqrt(14))/21 --------------------------------------- x=(14+i sqrt(14))/(21) , (14-i sqrt(14))/(21)

OpenStudy (amistre64):

straight for the juggular ....